OpenStudy (anonymous):

A square has sides length x cm. A new rectangle is created by increasing one dimension by 4 cm and decreasing the other dimension by 5 cm. What is the expanded and factored equations for area?

OpenStudy (anonymous):

if anyone could answer this that would be really great!

OpenStudy (anonymous):

L=x+4 W=x-5 Area= L*W =(x+4)(x-5) =x^2-x-9